can you solve the system of equations, y=5x-5 and y=1x+7
Answer by mananth(16946)

-5 x + 1 y = -5 .............1
-1 x + 1 y = 7 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -1
-5 x 1 y = -5
1 x -1 y = -7
Add the two equations
-4 x = -12
/ -4
x = 3
plug value of x in (1)
-5 x + 1 y = -5
-15 + 1 y = -5
1 y = -5 + 15
1 y = 10
y = 10